OpenStudy (anonymous):

If g(3) = 18, then g(x) = 2x + _____? is this the same??

OpenStudy (anonymous):

12

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

g(3) = 18 means that the point (3,18) is on the line

OpenStudy (anonymous):

and the g(x) is already in slope intercept form, so all you need to do is find the y intercept

OpenStudy (anonymous):

im still so confused :/

OpenStudy (anonymous):

you know y = mx + b?

OpenStudy (anonymous):

yes been forever since i done y intercept but yes i know

OpenStudy (anonymous):

g(x) is the same as f(x) is the same as y as far as notation goes don't let it confuse you

OpenStudy (anonymous):

we know a point because g(3) = 18 means that (3,18) is on the line

OpenStudy (anonymous):

and we have most of the equation already with the y = 2x + b (b means blank)

OpenStudy (anonymous):

so just plug in the (x=3, y=18) in y = 2x + b and solve for b

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

sorry i was writing it down so ur basically putting it in y intercept form?

OpenStudy (anonymous):

it already was

OpenStudy (anonymous):

18y=2(3)=-----

OpenStudy (anonymous):

+---- sorry im horrible at math

OpenStudy (anonymous):

thats almost it... but you replaced the y with 18, so the y went away 18 = 2(3) + b

OpenStudy (anonymous):

so be 18-6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks :)